# Video-Foley: Two-Stage Video-To-Sound Generation via Temporal Event Condition For Foley Sound

### v2

@@ Abstract
Foley sound synthesis is crucial for multimedia production, enhancing user experience by synchronizing audio and video both temporally and semantically. Recent studies on automating this labor-intensive process through video-to-sound generation face significant challenges. Systems lacking explicit temporal features suffer from poor alignment and controllability, while timestamp-based models require costly and subjective human annotation. We propose Video-Foley, a video-to-sound system using Root Mean Square (RMS) as an intuitive condition with semantic timbre prompts (audio or text). RMS, a frame-level intensity envelope closely related to audio semantics, acts as a temporal event feature to guide audio generation from video. The annotation-free self-supervised learning framework consists of two stages, Video2RMS and RMS2Sound, incorporating novel ideas including RMS discretization and RMS-ControlNet with a pretrained text-to-audio model. Our extensive evaluation shows that Video-Foley achieves state-of-the-art performance in audio-visual alignment and controllability for sound timing, intensity, timbre, and nuance. Source code, model weights and demos are available on our companion website [^1] https://jnwnlee.github.io/video-foley-demo.
Index Terms:
Video-to-Sound, Video-to-Audio, Controllable Audio Generation, Multimodal Deep Learning.
@@ I Introduction
We propose Video-Foley, a two-stage model that leverages temporal event conditions for annotation-free training of highly synchronized Foley sound generation. At its core, we introduce the Root Mean Square (RMS) of audio content as a key temporal feature for audio-visual synchrony. Defined as a frame-level energy feature calculated from audio waveforms, RMS captures not only the presence of sound events but also their intensity and temporal change, associated with subtle timbre and nuance. By incorporating RMS alongside audio or text prompts, our approach ensures high temporal and semantic synchronization, enhancing controllability. As illustrated in Fig. 1, the proposed two-stage framework first predicts RMS values from video frames (Video2RMS) and then generates temporally aligned audio waveforms based on the predicted RMS (RMS2Sound). Our model is trained efficiently with video-audio pairs (i.e., general video files) and audio-only data without human annotations, using novel techniques such as RMS discretization and RMS-ControlNet leveraging a pretrained text-to-audio model. Through objective evaluation metrics and subjective human listening test, we demonstrate that Video-Foley achieves state-of-the-art performance in both temporal and semantic alignment on the Greatest Hits dataset. Additionally, qualitative analysis and accompanying demo highlight its high controllability over timing, intensity, timbre, and nuance in the generated audio.
@@ II Related Works
Figure 2: Two model architecture types in neural video-to-sound generation: single-latent model (top) and dual-latent model (bottom).
@@ II-B Neural Audio Generation with Temporal Event Condition
Various neural audio generation approaches have leveraged these temporal features to enable temporal event control. A straightforward approach is to train a model from scratch that directly takes these conditioning inputs. T-Foley [14] introduced a waveform domain diffusion model conditioned on both RMS and sound category text to guide audio generation. Their results demonstrated that RMS effectively controls the temporal characteristics while also influencing semantic elements such as timbre. Moreover, T-Foley showed that RMS can be easily manipulated by users through simple actions like voice or clap sounds.
Other methods have focused on enhancing the controllability of pretrained large-scale generative models by incorporating new input types, leveraging these models’ performance and generalizability. In the vision domain, ControlNet [19] introduces additional spatial conditioning controls—such as sketch, depth, and human pose—into a high-performance text-to-image latent diffusion model (LDM). Instead of fine-tuning the entire U-Net-based diffusion model, ControlNet freezes the original model, duplicates its encoding layers, and finetunes the copies to learn additional control conditions. IP-Adapter[20] enhances controllability by introducing decoupled cross-attention layers, which process image features separately from text. This method finetunes fewer parameters and leverages a pretrained image encoder for feature extraction. In the audio domain, Guo et al. [21] proposed Fusion-Net, which incorporates fine-grained temporal inputs—such as timestamps, pitch contour, and energy contour (similar to RMS)—into a pretrained Text-to-Audio (TTA) model. Their method applies convolutions, linear projections, and self-attention layers. Music ControlNet [22] extends ControlNet to the music domain, enabling time-varying control in a pretrained text-to-music diffusion model through three distinct inputs: chromagram for melody, frame-wise energy for dynamics, and beat/downbeat logits for rhythm.
@@ III-A Overview
Figure 1 illustrates our proposed neural video-to-sound generation model, Video-Foley. It consists of two parts: Video2RMS and RMS2Sound. The overall pipeline operates as follows: first, the model takes a video and a prompt as inputs. The prompt, which describes the desired sound, can be either an audio sample or a text description, corresponding to the semantic timbre condition in Fig. 2. Next, Video2RMS predicts the temporal feature of the audio from video, capturing its temporal dynamics. The output serves as the temporal event condition defined in Fig. 2. Finally, RMS2Sound takes both the semantic and temporal conditions and generates the corresponding sound through a diffusion process.
@@ III-B RMS as a Temporal Event Condition
There are two main reasons for selecting RMS. First, timestamps capture only the start and end points of sound events, overlooking crucial audio characteristics such as volume dynamics—for instance, the gradual intensity shift of a moving car, which is difficult to describe in text. RMS effectively represents these aspects, from transient event-based sounds to continuous ambient sounds [14]. Additionally, human annotation for timestamp is costly and highly subjective, as further discussed in the next paragraph. Moreover, Heller et al. [24] demonstrated that listeners perceive hybrid sounds—Foley sounds combined with the temporal envelope of real recordings—as more realistic than either Foley or real recordings alone. This finding underscores the importance of intensity dynamics in sound generation.
@@ III-C Video2RMS
As illustrated in Fig. 4, the Video2RMS model consists of three 1D-convolutional blocks, two Bi-LSTM layers, and a linear projection head, similar to the visual encoder of RegNet [1]. For input, the BN-Inception network [30], pretrained on ImageNet [^2] https://yjxiong.blob.core.windows.net/models/bn_inception-9f5701afb96c8044.pth, extracts video features frame-wise from RGB images and 2-channel optical flows. For optical flow extraction, pretrained RAFT (Raft_Large_Weights.C_T_SKHT_V2) in pytorch was used.333 https://pytorch.org/vision/main/models/generated/torchvision.models.optical_flow.raft_large.html The two features are then concatenated for each time frame. Three convolutional blocks process the local information of the feature sequence. Each convolutional block includes a convolution layer, a batch normalization layer, and a ReLU activation layer. Two layers of bidirectional LSTM encode the global information of the features across the time axis. Finally, the linear head projects the feature sequence to predict the classification probability for each RMS bin. The loss function is defined as $L=\sum_{i}CE(\hat{c}_{i},c_{i})$ where $c_{i}$ denotes the discretized RMS class label at i-th frame, $\hat{\cdot}$ is the prediction, and $CE$ is the cross entropy loss.
@@ III-D RMS2Sound
To guide the audio generation that reflects both semantic and temporal conditions, we propose RMS2Sound which is a combination of RMS-ControlNet and a frozen TTA model, that generates audio from input RMS and audio-text joint embedding as shown in Fig. 4. RMS-ControlNet incorporates additional RMS conditions into a pretrained TTA model, i.e., AudioLDM [6]. Leveraging CLAP[31] embedding as a joint audio-text space, RMS2Sound can generate audio from either text or audio prompts. We train RMS-ControlNet on audio-only data to leverage its larger scale compared to video datasets. The training loss function is as follows:
@@
where $\epsilon$ is the noise injected during forward diffusion process, $x$ is the audio waveform, $z$ is a latent representation of $x$ encoded with a pretrained variational autoencoder, $z_{t}$ is $z$ at $t$ diffusion timestep, $C$ is the CLAP encoder, and $R$ is the RMS computation. RMS-ControlNet has a similar architecture to the original ControlNet [19]. We freeze the parameters of the TTA model and train only a copy of its encoding layers. The zero-initialized convolution layers connect ControlNet to the pretrained network for stable training. To match the feature size of RMS to that of $z_{t}$, we use an additional 2D zero convolutional layer.
@@ IV-A Dataset
We used the Greatest Hits dataset[26] with its official train-test split for training and evaluation. The dataset contains 977 videos of a person making sounds with a wooden drumstick on 17 different materials (wood, metal, rock, leaf, plastic, cloth, water, etc.) using two types of actions (hit, scratch). We segmented the videos with denoised audio into 10-second clips without overlap, and resampled to 16kHz for audio and 30fps for video. Each video frame was resized to 344$\times$256 pixels. This resulted in 2,212 training videos (6.14 hours) and 732 test videos (2.03 hours). The training set was used to train Video2RMS, and the test set was used to evaluate both Video2RMS and the entire Video-Foley model. To increase extensibility and applicability, we trained RMS-ControlNet using audio-only data from a variety of sounds, rather than limiting it to hit and scratch sounds. We used the FreeSound dataset[32], which contains about 6,000 hours of audio. All audio was resampled to 16 kHz.
@@ IV-B1 Training
The Video2RMS and RMS2Sound models are trained separately but combined during inference. For Video2RMS, RMS was calculated from the audio waveform with a 512 window size and a 128 hop length, following the configuration in T-Foley[14]. By padding $(512-128)/2$ values at both ends of the waveform in reflect mode, we obtained 1250 frames. Then, the RMS was discretized into 64 bins ($\simeq$0.5dB granularity), and Gaussian label smoothing was applied ($W=2)$. The model was trained for 500 epochs using a StepLR scheduler (rate $1e$-$3$, step size 100), with a batch size of 512 using Adam optimizer. For RMS2Sound, the window and hop length of RMS are set to 1024 and 160 respectively, following AudioLDM’s detail. By padding $(1024-160)/2$ values at both ends of the waveform in reflect mode, we obtained 1024 frames. When using the predicted RMS from Video2RMS, nearest-neighbor interpolation is applied to match the feature length. Note that the RMS was not discretized in RMS-ControlNet, i.e. the continuous-valued RMS is the conditioning input. AudioLDM was used as the backbone TTA model. We initialized the weights of both AudioLDM and ControlNet using an official checkpoint ‘audioldm-s-full’. RMS-ControlNet was trained for 300k steps using the AdamW optimizer. To maintain training consistency, we adhere to the original AudioLDM configuration (e.g., audio normalization). The learning rate started at $1e$-$4$ and was halved every 10k steps. During training, only the parameters of RMS-ControlNet were updated.
@@ IV-C Baseline Models
For comparison with our model, 6 different video-to-sound generation models were chosen. SpecVQGAN [8] and Diff-Foley [10] correspond to single-latent approaches, while CondFoleyGen [17] with audio-visual prompt and SyncFusion [12] utilizing onset fall into dual-latent models along with ours, described in Section II-A. Although some models (i.e., SpecVQGAN and Diff-Foley) do not receive text or audio as semantic prompts, we included them in the evaluation to compare the performance mainly on temporal alignment. Regarding the training dataset, SpecVQGAN was trained on VGGSound[33] ($\sim$0.4k hours), Diff-Foley was trained on VGGSound and AudioSet[34] ($\sim$1.3k hours), which are in-the-wild datasets based on Youtube videos, larger than Greatest Hits. CondFoleyGen and SyncFusion were trained on Greatest Hits ($\sim$6 hours). We used official code and checkpoints for every model. Unless otherwise noted, we followed the default configuration choices for inference. SpecVQGAN generates 10 seconds of audio from 21.5 fps video. The model ‘2021-07-30T21-34-25_vggsound_transformer’ was used. As Diff-Foley generates 8-second audio from 4 fps video, we made two inferences: one with video frames from 0-8 seconds and another from 2-10 seconds. We then concatenated the entire first segment with the latter 2 seconds of the second segment to produce a 10-second audio. Although CondFoleyGen generates 2 seconds of audio from 15 fps video, the official code was implemented to generate multiples of 2 seconds of audio by adjusting the parameter $W_{scale}$. We set $W_{scale}$ to 5 to generate 10 seconds of audio. SyncFusion was trained to generate 5.46 seconds of audio from 15fps video. We generated 5-second audio clips and concatenated them. For text prompts, we used the same text as Video-Foley.
For Video2RMS prediction, we provide the model trained in a regression setting, mentioned in Section III-C, as a baseline. This model is trained with L2 loss $L=\sum_{i}||\hat{R}_{i}-R_{i}||^{2}$ where $R_{i}$ denotes the continuous RMS value of i-th frame, $\hat{\cdot}$ is the prediction. Other details, such as model architecture or configurations, are the same as in our proposed classification model.
@@ IV-D1 Objective Evaluation
Additionally, the CLAP [41] score was calculated by measuring the cosine distance between the generated and ground-truth audio pairs in the joint text-audio embedding space. [^5] Note that this CLAP model is from LAION [41], which differs from the Microsoft model [31] used in AudioLDM. First, we extract embeddings from ground-truth $e$ and generated audio $\hat{e}$ in the audio-text joint embedding space of CLAP. Then, the cosine distance between the two embedding vectors $\cos(e,\hat{e})$ is measured.
$E\text{-}L1={1\over k}\displaystyle\Sigma_{i=1}^{k}||E_{i}-\hat{E}_{i}||$ (7)
where $E_{i}$ is the ground-truth event feature of $i$-th frame, and $\hat{E}_{i}$ is the predicted one. In this paper, RMS is the temporal event feature. For evaluating Video2RMS, E-L1 between the predicted RMS and the ground-truth RMS is measured. In the case of Video-Foley, E-L1 between the RMS extracted from generated audio and ground-truth audio are considered.
All metrics except FAVD were used to evaluate RMS2Sound, as there is no video input. Along with E-L1, classification accuracy with tolerance windows of about $\pm$1/3/6 dB ($\pm$2/5/8 adjacent class bins) measured the RMS prediction performance of Video2RMS, excluding frames where both ground truth and prediction are silent, similar to the previous study [17]. This exclusion is necessary because only a small portion of audio frames are non-silent for hit/scratch actions, making the model learn an undesired shortcut for predicting silence and thus failing to effectively capture the performance in non-silent frames.
For some ablation studies, Onset metrics, Onset Acc (accuracy) and Onset AP (average precision), were measured following the protocol in CondFoleyGen[17]. These metrics were calculated with a tolerance of $\pm$0.1 seconds after applying non-maximum suppression with a window length of 50 ms based on the onset prediction confidence. Moreover, these metrics also exclude true-negative cases (both predicted and labeled outputs being non-onset) due to the sparsity of onset labels.
@@
random choice (l.b.) 0.2807 0.077 0.165 0.248
Regression (baseline) 0.0192 0.126 0.229 0.285
Classification (Ours) 0.0150 0.164 0.349 0.498
     w/ label smoothing 0.0147 0.165 0.361 0.506
disc. RMS (g.t.) 0.00243 1.000 1.000 1.000
Figure 6: Comparison of the ground-truth RMS curve (left) with the predicted curves from the regression baseline (middle) and our classification model (right).

### v3

@@ Abstract
Foley sound synthesis is crucial for multimedia production, enhancing user experience by synchronizing audio and video both temporally and semantically. Recent studies on automating this labor-intensive process through video-to-sound generation face significant challenges. Systems lacking explicit temporal features suffer from poor alignment and controllability, while timestamp-based models require costly and subjective human annotation. We propose Video-Foley, a video-to-sound system using Root Mean Square (RMS) as an intuitive condition with semantic timbre prompts (audio or text). RMS, a frame-level intensity envelope closely related to audio semantics, acts as a temporal event feature to guide audio generation from video. The annotation-free self-supervised learning framework consists of two stages, Video2RMS and RMS2Sound, incorporating mu-law scaled RMS discretization and RMS-ControlNet with a pretrained text-to-audio model. Our extensive evaluation shows that Video-Foley achieves state-of-the-art performance in audio-visual alignment and controllability for sound timing, intensity, timbre, and nuance. Source code, model weights and demos are available on our companion website [^1] https://jnwnlee.github.io/video-foley-demo.
@@ I Introduction
We propose Video-Foley, a two-stage model that leverages temporal event conditions for annotation-free training of highly synchronized Foley sound generation. Rather than generating audio directly from video frames, our model first predicts a temporal feature as an intermediate representation, then generates audio from it. At its core, we introduce the Root Mean Square (RMS) of audio content as a key temporal feature. Defined as a frame-level energy feature calculated from audio waveforms, RMS captures not only the presence of sound events but also their intensity and temporal change, associated with subtle timbre and nuance [14, 15]. We propose to incorporate RMS as a target in the video encoding stage to ensure strong temporal and semantic audio-visual synchronization. Together with audio or text prompts, RMS serves as a control condition in the audio generation stage, enhancing controllability. Our two-stage framework, illustrated in Fig. 1, consists of Video2RMS and RMS2Sound. Video2RMS first predicts the RMS curve from video effectively using techniques such as label discretization and smoothing. Subsequently, RMS2Sound takes the RMS with an audio or text prompt to generate a temporally and semantically aligned audio waveform. Inspired by ControlNet [16], designed to add spatial conditioning (e.g., sketch) to pretrained image generators, our proposed RMS-ControlNet guides a frozen text-to-audio model [6] via an RMS curve. The two modules are trained separately: Video2RMS trained on video-audio pairs (i.e., general video files), while RMS2Sound trained on audio-only data—both without any human annotations. Through objective evaluation metrics and subjective human listening test, we demonstrate that Video-Foley achieves state-of-the-art performance in both temporal and semantic alignment on the Greatest Hits dataset [17]. Additionally, qualitative analysis and accompanying demo highlight its high controllability over timing, intensity, timbre, and nuance in the generated audio.
@@ II Related Works
Figure 2: Two model architecture types in neural video-to-sound generation: (a) single-latent model and (b) dual-latent model.
@@ II-B Neural Audio Generation with Temporal Event Condition
Various neural audio generation approaches have leveraged these temporal features to enable temporal event control. A straightforward approach is to train a model from scratch that directly takes these conditioning inputs. T-Foley [14] introduced a waveform domain diffusion model conditioned on both RMS and sound category text to guide audio generation. The results from T-Foley and other follow-up studies, such as MambaFoley [15], demonstrated that RMS effectively controls the temporal characteristics while also influencing semantic elements such as timbre. Moreover, they showed that RMS can be easily manipulated by users through simple actions like voice or clap sounds.
Other methods have focused on enhancing the controllability of pretrained large-scale generative models by incorporating new input types, leveraging these models’ performance and generalizability. In the vision domain, ControlNet [16] introduces additional spatial conditioning controls—such as sketch, depth, and human pose—into a high-performance text-to-image latent diffusion model (LDM). Instead of fine-tuning the entire U-Net-based diffusion model, ControlNet freezes the original model, duplicates its encoding layers, and finetunes the copies to learn additional control conditions. IP-Adapter[22] enhances controllability by introducing decoupled cross-attention layers, which process image features separately from text. This method finetunes fewer parameters and leverages a pretrained image encoder for feature extraction. In the audio domain, Music ControlNet [23] extends ControlNet to the music domain, enabling time-varying control in a pretrained text-to-music diffusion model through three distinct inputs: chromagram for melody, frame-wise energy for dynamics, and beat/downbeat logits for rhythm. Guo et al. [24] proposed Fusion-Net, which incorporates fine-grained temporal inputs—such as timestamps, pitch contour, and energy contour (similar to RMS)—into a pretrained Text-to-Audio (TTA) model. Their method applies convolutions, linear projections, and self-attention layers.
@@ III-A Overview
Fig. 1 illustrates our proposed neural video-to-sound generation model, Video-Foley. It consists of two parts: Video2RMS and RMS2Sound. The overall pipeline operates as follows: first, the model takes a video and a prompt as inputs. The prompt, which describes the desired sound, can be either an audio sample or a text description, corresponding to the semantic timbre condition in Fig. 2. Next, Video2RMS predicts the temporal feature of the audio from video, capturing its temporal dynamics. The output serves as the temporal event condition defined in Fig. 2. Finally, RMS2Sound takes both the semantic and temporal conditions and generates the corresponding sound through a diffusion process.
@@ III-B RMS as a Temporal Event Condition
There are two main reasons for selecting RMS. First, timestamps capture only the start and end points of sound events, overlooking crucial audio characteristics such as volume dynamics—for instance, the gradual intensity shift of a moving car, which is difficult to describe in text. RMS effectively represents these aspects, from transient event-based sounds to continuous ambient sounds [14], and has already been shown to serve as an effective temporal condition for audio generation, as discussed in Section II-B. Additionally, human annotation for timestamp is costly and highly subjective, as further discussed in the next paragraph. Moreover, Heller et al. [26] demonstrated that listeners perceive hybrid sounds—Foley sounds combined with the temporal envelope of real recordings—as more realistic than either Foley or real recordings alone. This finding underscores the importance of intensity dynamics in sound generation.
@@ III-C Video2RMS
As illustrated in Fig. 4, the Video2RMS model consists of three 1D-convolutional blocks, two Bi-LSTM layers, and a linear projection head. The architecture is inspired by the visual encoder of RegNet [1], with the key difference that our model includes a linear head to predict the RMS curve using a classification loss, whereas RegNet uses the LSTM output as a temporal video feature implicitly trained with GAN loss. For input, the BN-Inception network [31], pretrained on ImageNet classification [^2] https://yjxiong.blob.core.windows.net/models/bn_inception-9f5701afb96c8044.pth, extracts video features frame-wise from RGB images and 2-channel optical flows. For optical flow extraction, pretrained RAFT (Raft_Large_Weights.C_T_SKHT_V2) in pytorch was used. [^3] https://pytorch.org/vision/main/models/generated/torchvision.models.optical_flow.raft_large.html Since BN-Inception is originally designed for 3-channel image inputs, the first convolutional kernel is inflated by averaging across the three channels and duplicating across two axes to accommodate the two-channel optical flow. The feature is taken after the last average pooling layer of the frozen BN-Inception. The two features are then concatenated for each time frame. Three convolutional blocks process the local information of the feature sequence. Each convolutional block includes a convolution layer, a batch normalization layer, and a ReLU activation layer. Two layers of bidirectional LSTM encode the global information of the features across the time axis. Finally, the linear head projects the feature sequence to predict the classification probability for each RMS bin. The loss function is defined as $L=\sum_{i}CE(\hat{c}_{i},c_{i})$ where $c_{i}$ denotes the discretized RMS class label at i-th frame, $\hat{\cdot}$ is the prediction, and $CE$ is the cross entropy loss.
@@ III-D RMS2Sound
To guide the audio generation that reflects both semantic and temporal conditions, we propose RMS2Sound which is a combination of RMS-ControlNet and a frozen TTA model, that generates audio from input RMS and audio-text joint embedding as shown in Fig. 4. RMS-ControlNet consists of a trainable copy of the encoding layers and the middle block of the backbone TTA model, connected to the frozen backbone layer-wise through zero-initialized convolutional layers. AudioLDM [6], conditioned on CLAP [32] embeddings, was used as the backbone TTA model. RMS-ControlNet receives the same input as AudioLDM, except that the noisy latent is summed with the RMS condition. To match the feature dimensions of the RMS condition to those of the noisy latent, we apply a 2D zero-initialized convolutional layer. RMS-ControlNet is trained following the same procedure as the original ControlNet [16]. The training loss function is as follows:
@@
where $\epsilon$ is the noise injected during forward diffusion process, $x$ is the audio waveform, $z$ is a latent representation of $x$ encoded with a pretrained variational autoencoder, $z_{t}$ is $z$ at $t$ diffusion timestep, $C$ is the CLAP encoder, and $R$ is the RMS computation. We freeze the parameters of AudioLDM and update only those of RMS-ControlNet. RMS-ControlNet is trained on audio-only data to take advantage of its larger scale compared to video datasets. Since CLAP provides a joint audio-text representation space, RMS2Sound is capable of generating audio from either text or audio prompts.
@@ IV-A Dataset
[…]
@@ IV-B1 Training
The Video2RMS and RMS2Sound models are trained separately but combined during inference. For Video2RMS, RMS was calculated from the audio waveform with a 512 window size and a 128 hop length, following the configuration in T-Foley[14]. By padding (512128)/2(512-128)/2 values at both ends of the waveform in reflect mode, we obtained 1250 frames. Then, the RMS was discretized into 64 bins ($\simeq$0.5dB granularity), and Gaussian label smoothing was applied ($W=2)$. The model was trained for 500 epochs using a StepLR scheduler (rate 1e1e-33, step size 100), with a batch size of 512 using Adam optimizer. For RMS2Sound, the window and hop length of RMS are set to 1024 and 160 respectively, following AudioLDM’s detail. By padding (1024160)/2(1024-160)/2 values at both ends of the waveform in reflect mode, we obtained 1024 frames. When using the predicted RMS from Video2RMS, nearest-neighbor interpolation is applied to match the feature length. Note that the RMS was not discretized in RMS-ControlNet, i.e. the continuous-valued RMS is the conditioning input. We initialized AudioLDM using the official checkpoint ‘audioldm-s-full’ [^4] https://huggingface.co/haoheliu/AudioLDM-S-Full. For ControlNet, we used only the weights of the encoder and middle block of the U-Net in the same checkpoint. RMS-ControlNet was trained for 300k steps using the AdamW optimizer. To maintain training consistency, we adhere to the original AudioLDM configuration (e.g., audio normalization). The learning rate started at 1e1e-44 and was halved every 10k steps. During training, only the parameters of RMS-ControlNet were updated.
@@ IV-C Baseline Models
For comparison with our model, we include two primary baselines for video-to-sound generation: CondFoleyGen [20], which uses audio-visual prompt, and SyncFusion [12], which leverages onset timestamps. Both, like our model, fall under the dual-latent category described in Section II-A. These baselines help validate our choice of RMS as a bridging temporal feature. All models were trained on Greatest Hits ($\sim$6 hours). We also include additional single-latent baselines, SpecVQGAN [8] and Diff-Foley [10]. Note that these models do not receive any text or audio as semantic prompts, putting them at a disadvantage in terms of semantic alignment. SpecVQGAN was trained on VGGSound[34] ($\sim$0.4k hours), Diff-Foley was trained on VGGSound and a subset of AudioSet[35] ($\sim$1.1k hours). Despite the larger size, these in-the-wild datasets sourced from Youtube suffer from low audio-visual quality, duplicates, and visually non-indicative sounds (i.e., weak alignment between visual content and sound such as off-screen sound or background noise [36, 37]). While acknowledging that a perfectly fair comparison is difficult, we include these representative single-latent models to provide broader context in both objective and subjective evaluation.
All baseline inferences were conducted using official code and checkpoints. Unless otherwise noted, we followed the default configuration choices for inference. Although CondFoleyGen generates 2 seconds of audio from 15 fps video, the official code was implemented to generate multiples of 2 seconds of audio by adjusting the parameter $W_{scale}$. We set $W_{scale}$ to 5 to generate 10 seconds of audio. SyncFusion was trained to generate 5.46 seconds of audio from 15fps video. We generated 5-second audio clips and concatenated them. For text prompts, we used the same text as Video-Foley. SpecVQGAN generates 10 seconds of audio from 21.5 fps video. The model ‘2021-07-30T21-34-25_vggsound_transformer’ was used. As Diff-Foley generates 8-second audio from 4 fps video, we made two inferences: one with video frames from 0-8 seconds and another from 2-10 seconds. We then concatenated the entire first segment with the latter 2 seconds of the second segment to produce a 10-second audio.
[…]
@@ IV-D1 Objective Evaluation
random choice (l.b.) 0.299 0.077 0.165 0.248
Regression (baseline) 0.119 0.126 0.229 0.285
Classification (Ours) 0.082 0.164 0.349 0.498
   w/ label smoothing 0.080 0.165 0.361 0.506
disc. RMS (g.t.) 0.018 1.000 1.000 1.000
[…]
where $E_{i}$ is the ground-truth event feature of $i$-th frame, and $\hat{E}_{i}$ is the predicted one. In this paper, RMS scaled with $\mu$-law encoding is the temporal event feature. For evaluating Video2RMS, E-L1 between the predicted RMS and the ground-truth RMS is measured. In the case of Video-Foley, E-L1 between the RMS extracted from generated audio and ground-truth audio are considered.
Figure 6: Comparison of the ground-truth RMS curve (a) with the predicted curves from the regression baseline (b) and our classification model (c).
Figure 7: Performance of Video2RMS (a) and RMS2Sound (b) for different numbers of RMS bins. w/o_rms: without RMS condition (Text-to-Audio[6]), cont.: continuous RMS, no discretization.
All metrics except FAVD were used to evaluate RMS2Sound, as there is no video input. Classification accuracy with tolerance windows of about $\pm$1/3/6 dB ($\pm$2/5/8 adjacent class bins) measured the RMS prediction performance of Video2RMS, excluding frames where both ground truth and prediction are silent, similar to the previous study [20]. This exclusion is necessary because only a small portion of audio frames are non-silent for hit/scratch actions, making the model learn an undesired shortcut for predicting silence and thus failing to effectively capture the performance in non-silent frames.